OpenStudy (anonymous):

simplify: -1/4 divided by 5

OpenStudy (tkhunny):

Further simplify: \(\dfrac{-1}{4}\cdot \dfrac{1}{5}\)

OpenStudy (anonymous):

is it -1/20 ?

OpenStudy (anonymous):

-1/4.5/1=-5/4 or -1.25 it solv ok

OpenStudy (tkhunny):

@JennWoww Please get a little more confidence than that. Dividing fractions required "reciprocal and multiply". Did we do that? Did we do it right? Why do you doubt? Multiplying fractions requires "multiply numerators and multiply denominators". Did we do that? Did we do it right? Why do you doubt? You tell me if it is correct.

OpenStudy (anonymous):

ok i am try agian 

OpenStudy (tkhunny):

Please avoid the extensive decimal places. Exact valeus from fractions are most often more suitable.